### v3

# Feasibility of Learning Weighted Automata on a Semiring

Abstract.
Since the seminal work by Angluin and the introduction of the $L^{*}$-algorithm, active learning of automata by membership and equivalence queries has been extensively studied to learn various extensions of automata. For weighted automata, algorithms for restricted cases have been developed in the literature, but so far there was no global approach or understanding how these algorithms could (or could not) apply in the general case. In this paper we chart the boundaries of the Angluin approach (using a class of hypothesis automata constructed from membership and equivalence queries) for learning weighted automata over a general semiring. We show precisely the theoretical limitations of this approach and classify functions with respect to how ‘guessable’ they are (corresponding to the existence and abundance of solutions of certain systems of equations). We provide a syntactic description of the boundary condition for a correct hypothesis of the prescribed form to exist. Of course, from an algorithmic standpoint, knowing that (many) solutions exist need not translate into an effective algorithm to find one; we conclude by relating our work to the existing literature with a discussion of some known conditions (and variants thereof) that suffice to ensure this, illustrating the ideas over several familiar semirings (including the natural numbers).
Key words and phrases:
Weighted Automata, Learning, Semiring, Angluin algorithm
1. Introduction
Imagine the following situation: you want to model a computer system (for verification purposes for example) with some mathematical abstraction, but the internal state of the system cannot be accessed. Think for example of very complex systems such as AI systems, that can realistically only be viewed as black-boxes: the system can be tested, you can feed it some chosen inputs and observe the outputs. Given these (finitely many) input-output pairs, you create a model. Suppose you also have a way to decide whether your model matches (or, more realistically in the application fields, is close enough to) the system’s behaviour. If this is not the case, the system gives you an input that behaves incorrectly in your model and using this new input-output pair together with further observations, you construct a new model to test. This is active learning.
Automata learning:
In her seminal paper [4], Angluin introduced the $L^{*}$-algorithm where a learner tries to guess a rational language $L$ known only by an oracle. The learner is allowed to ask two types of queries: (1) membership queries where the learner chooses a word $w$ and asks the oracle whether $w$ belongs to $L$, and (2) equivalence queries where the learner chooses an automaton $\mathcal{A}$ and asks the oracle whether $\mathcal{A}$ recognises $L$; if this is not the case, the oracle gives a word (called counter-example) witnessing this fact. The $L^{*}$-algorithm allows the learner to correctly guess the minimal deterministic automaton recognising $L$ in a number of membership and equivalence queries polynomial in its size and the length of counter-examples given by the oracle. An overview of applications of automata learning can be found in [27, 24, 35]. In particular, the $L^{*}$-algorithm has been used in [40] to extract automata from recurrent neural networks.
Weighted automata:
Weighted automata are a quantitative extension of automata where transitions are weighted in a semiring, allowing to model probabilities, costs or running time of programs [34, 17, 18]. They find applications in image compression [25, 14], language and speech processing [30, 31], bioinformatics [1], formal verification [12, 3], analysis of on-line algorithms [2] and probabilistic systems [36]. Notably, they have also been used recently to model recurrent neural networks [33, 32, 41, 39].
Extending automata learning:
Active learning has been extended to weighted automata but only for restricted classes: it has first been studied for fields [8, 7], and then principal ideal domains [21]. The Angluin framework has also been generalised to other extensions of automata and transducers [37, 38, 11, 9, 6, 5, 10, 16, 29, 23]. In [13] (and at first [22]), a categorical approach was proposed to encompass many of these cases. However, although this encompasses the case of weighted automata over fields, this framework does not include weighted automata in general. As often with weighted automata, finding a unified approach (working for all semirings) is not easy.
In [21], van Heerdt, Kupke and Silva set a general framework for learning weighted automata, presenting a general learning algorithm, and focus on specifying two conditions on the semirings (ascending chain condition and progress measure), ensuring termination of the algorithm - this allows them to develop a learning algorithm à la Angluin for automata weighted on a principal ideal domain. Section 5 and Theorem 17 of [21] also present a specific example of an automaton (as well as a more general class) over the non-negative integers for which their algorithm does not terminate. In the present paper, we develop a general framework that gives a deeper understanding of the reasons for that.
Contribution:
Our approach is somewhat different to what has been done so far: we look globally at all functions computed by finite state weighted automata and investigate whether they can potentially be learned by an Angluin-style algorithm. Working over an arbitrary semiring, we define levels of learnability and a hierarchy of functions with respect to these levels: the ones that definitely cannot be learned à la Angluin, the ones that might be, provided we are lucky or have a good algorithmic strategy, and the ones that definitely will be. Our main result in this regard is a syntactic condition on the functions that have the potential to be learned by an Angluin-style algorithm. For such a function there must exist a finite state automaton of a prescribed type computing it: we call these automata literal and they are closely related to residual automata. In cases already widely studied in the literature (e.g. where the weights are drawn from fields or integral domains), it turns out that every function computed by a finite state automaton can be computed by a literal automaton. Thus, in the existing literature, focus is (reasonably) on algorithmic issues. In this paper, we focus on picturing the landscape for active learning in arbitrary semirings, and showing the limitations; for instance over the non-negative integers, it is easily seen that the example of Heerdt, Kupke and Silva mentioned in the previous paragraph cannot be computed by a literal automaton. We do discuss a general algorithm (the one from [21]) and give some conditions on the target functions to ensure termination. In doing so, we refine the ascending chain condition and progress measure from [21] to pinpoint exactly what is needed, and we explain fully the reasons behind [21, Theorem 17].
Within this framework, we then consider some properties of semirings that cause certain parts of our general hierarchy to collapse, and consider several specific well-studied examples of semirings (including the non-negative integers) in this context and provide a more detailed picture of the hierarchy for these. Note that the learning algorithm is based on equivalence queries, and one might ask whether the undecidability of equivalence for certain semirings [26] is problematic. Since in the application fields, equivalence can rarely be used as such (and often testing or approximation might be used instead), we believe having a framework that applies even in these cases is still of interest.
The difficulty of this work was in setting a new framework and the (we believe) right notions for it. Once the correct definitions are in place, the results and proofs follow nicely. This new framework also opens avenues for research.
Structure of the paper:
In Section 2, we recall the definition of weighted automata, some specific examples, and outline the type of learning we are considering. We end this section with a set of questions that arise naturally, and which motivate the definitions and results in the rest of the paper. In Section 3, we explain what we mean by levels of learnability and define classes of functions depending on them. We give several characterisations for these functions, and relate our work back to the existing literature (specifically,  [21]), which focuses on algorithmic issues in specific settings. In Section 4, we picture the hierarchy based on the previous definitions, prove its strictness and give conditions which cause parts of the hierarchy to collapse. Throughout we raise a number of questions and indicate avenues for future progress, summarising these in 5.
2. Learning weighted automata
In this section, we recall the definition of weighted automata (on a semiring) and the type of learning we are considering.
2.1. Weighted automata and specific semirings
A finite alphabet is a finite set of symbols, called letters. A finite word is a finite sequence of letters and $\varepsilon$ denotes the empty word. We will also denote by $|w|$ the length of a word $w$ and by $|w|_{a}$ the number of occurrences of the letter $a$ in $w$.
A monoid $(M,\ast,e)$ is a set $M$ equipped with a binary operation $\ast$ that is associative and has a neutral element $e$. A semiring $(S,\oplus,\otimes,0_{S},1_{S})$ is a set $S$ equipped with two operations such that $(S,\oplus,0_{S})$ and $(S,\otimes,1_{S})$ are monoids, $\oplus$ is commutative, $\otimes$ distributes on the left and the right of $\oplus$ and $0_{S}$ is a zero for $\otimes$. A semiring is said to be commutative if $\otimes$ is commutative. Given a semiring $(S,\oplus,\otimes,0_{S},1_{S})$, two matrices $A$ and $B$ over $S$ can be multiplied as usual by $(AB)_{i,j}=\bigoplus_{k=1,\ldots,n}(A_{i,k}\otimes B_{k,j})$, provided the number, $n$, of columns of $A$ is equal to the number of rows of $B$. It is straightforward to check that this multiplication is associative.
Definition 2.1.
A weighted automaton $\mathcal{A}$ on a semiring $(S,\oplus,\otimes,0_{S},1_{S})$ over alphabet $\Sigma$ is defined by: a finite set of states $Q$; initial-state vector $\alpha\in S^{Q}$ (which we view as a row vector); for all letters $a$ of $\Sigma$, transition matrices $M(a)\in S^{Q\times Q}$; and final-state vector $\eta\in S^{Q}$ (which we view as a column vector). The automaton $\mathcal{A}$ computes a function $f_{\mathcal{A}}:\Sigma^{*}\rightarrow S$ via $f_{\mathcal{A}}(w_{1}\cdots w_{k})=\alpha M(w_{1})\cdots M(w_{k})\eta$. We say that $f_{\mathcal{A}}(w)$ is the value computed by $\mathcal{A}$ on input $w$.
A weighted automaton can be seen as a (labelled, directed) graph with set of vertices $Q$, such that (i) for each state $p\in Q$ there is an initial weight $\alpha_{p}$ and a final weight $\eta_{p}$; and (ii) for each pair of states $p,q\in Q$ and each $a\in\Sigma$ there is a transition (labelled edge) $p\xrightarrow{a\,:\,m}q$ where $m=M(a)_{p,q}$ is called the weight of the transition (and if $m=0_{S}$, then we often omit this edge from the graph). A run from a state $p$ to a state $q$ is a (directed) path in the graph, its weight is the product of $\alpha_{p}$ by the product of the weights on the transitions (taken in the order traversed) followed by $\eta_{q}$. The value computed on a word $w$ is the sum of all weights of runs labelled by $w$. A state $p$ such that $\alpha_{p}$ (resp. $\eta_{p}$) is not $0_{S}$ might be referred to as an initial (resp. final) state. We use both the matrix and graph vocabulary in this paper, as convenient.
Throughout the paper, $\Sigma$ denotes a finite alphabet and $S$ a semiring.
Specific semirings and examples:
Besides giving a general framework valid for any semiring, we consider several specific well-studied semirings that arise frequently in application areas. We will consider the semiring (,+,×$(\mathbb{R},+,\times,0,1)$ and its restriction to non-negative reals $\mathbb{R}_{\geq 0}$, integers $\mathbb{Z}$, and non-negative integers $\mathbb{N}$. We also consider the semiring ({}$(\mathbb{R}\cup\{-\infty\},\max,+,-\infty,0)$ and its restrictions to ${}\mathbb{Z}\cup\{-\infty\}$ and ${}\mathbb{N}\cup\{-\infty\}$, denoted by $\mathbb{R}_{\max}$, $\mathbb{Z}_{\max}$ and $\mathbb{N}_{\max}$ and the Boolean semiring $\mathbb{B}$. Note that all these semirings are commutative. Finally, we will consider the non-commutative semiring with domain the finite subsets of $\Sigma^{*}$, with operations given by union and concatenation ($XY=\{xy\mid x\in X,y\in Y\}$) and neutral elements $\emptyset$ and $\{\varepsilon\}$. This semiring will be denoted $\mathcal{P}_{\rm fin}(\Sigma^{*})$.
Example 2.2.
We consider $\Sigma=\{a,b\}$. Figure 1 depicts several weighted automata where the initial (resp. final) states with weight $1_{S}$ are indicated with an ingoing (resp. outgoing) arrow, the other states have initial or final weights $0_{S}$. All six examples will be used in Section 4 and the example in Figure 1(a) will be used as a running example to illustrate the different notions we introduce. The automaton in Figure 1(a) is considered over $\mathbb{N}_{\max}$ and computes the number of $a$’s if the word starts with an $a$, the number of $b$’s if the word starts with a $b$, and $\infty$- ∞ on the empty word. The automaton depicted in Figure 1(b) will be considered over $\mathbb{R}_{\max}$ and computes the length of the longest block of consecutive $a$’s in a word. The automaton in Figure 1(c) is considered over $\mathbb{N}$ and on input $w$ computes: $2^{|w|_{a}}$ if $w$ starts with an $a$; $2^{|w|_{b}}$ if $w$ starts with a $b$; and 00 if $w$ is the empty word. The automaton from Figure 1(d) is viewed over $\mathbb{R}_{\geq 0}$ and computes $12^{n}-$ 1 on the words $a^{n}$ for all positive integers $n$, and 00 on any other word. Finally the automata in Figures 1(e) and 1(f) are over $\mathcal{P}_{\rm fin}(\Sigma^{*})$. The first one has outputs: $\emptyset$ on the empty word; $\{a^{|w|_{a}}\}$ on all words $w$ starting with an $a$; and $\{b^{|w|_{b}}\}$ on all words $w$ starting with a $b$. The second one outputs $\{a^{|w|_{a}},b^{|w|_{b}}\}$ on all input words $w$. It is worth mentioning here that several of the examples above could be viewed as examples over smaller semirings (e.g. Figure 1(b) could be considered over the semiring $\mathbb{N}_{\rm max}$. However, as we shall see below, changing the semiring over which one views a given example necessarily changes the space of hypothesis automata that one can construct using membership and equivalence queries.
(a) Over $\mathbb{N}_{\max}$: computes $|w|_{x}$ if $w$ starts with letter $x\in\{a,b\}$ and $\infty$- ∞ on $\varepsilon$.
(b) Over $\mathbb{R}_{\max}$: computes the length of the longest block of consecutive $a$’s in $w$.
(c) Over $\mathbb{N}$: computes $2^{|w|_{x}}$ if $w$ starts with letter $x\in\{a,b\}$ and 00 if $w=\varepsilon$.
(d) Over $\mathbb{R}_{\geq 0}$: computes $12^{n}-$ 1 if $w=a^{n}$ for some $n\geq 1$, and 00 otherwise.
(e) Over $\mathcal{P}_{\rm fin}(\Sigma^{*})$: computes $\{x^{|w|_{x}}\}$ if $w$ starts with letter $x\in\{a,b\}$ and $\emptyset$ if $w=\varepsilon$.
(f) Over $\mathcal{P}_{\rm fin}(\Sigma^{*})$: computes $\{a^{|w|_{a}},b^{|w|_{b}}\}$
Figure 1. Example weighted automata and the values computed on input word $w$.
Linear combinations:
Since we are going to consider non-commutative semirings, care needs to be taken when talking about linear combinations. Given a set $Y$, an element $x$ of $S^{Y}$ and $\lambda$ in $S$, $\lambda\otimes x$ denotes the element of $S^{Y}$ computed from $x$ by multiplying every component of $x$ by $\lambda$ on the left. Given a subset $X\subseteq S^{Y}$, a left-linear combination over $X$ is one of the form $\bigoplus_{x\in X}\lambda_{x}\otimes x$ for some $\lambda_{x}$ in $S$, with only finitely many $\lambda_{x}$ different from $0_{S}$. The left-semimodule generated by $X$ is then the subset of $S^{Y}$ containing those elements that can be written as a left-linear combination over $X$. An element of the left-semimodule generated by $X$ will simply be said to be left-generated by $X$.
Mirror:
Finally, we will use the mirror operation on functions $\Sigma^{*}\to S$. The mirror of a word $w=w_{1}w_{2}\cdots w_{n}$ where for all $i=1,\ldots,n$, $w_{i}\in\Sigma$ is the word $\bar{w}=w_{n}\cdots w_{2}w_{1}$. For $f:\Sigma^{*}\to S$, the mirror of $f$, denoted by $\bar{f}:\Sigma^{*}\to S$ is defined as $\bar{f}(w)=f(\bar{w})$. Given a weighted automaton $\mathcal{A}$, the mirror of $\mathcal{A}$ is obtained by reversing its transitions and swapping the initial-state vector and the final-state vector. Note that if $S$ is commutative and $f$ is computed by a weighted automaton $\mathcal{A}$ over $S$ then $\bar{f}$ is computed by the mirror of $\mathcal{A}$. However, there are cases when $S$ is non-commutative where a function can be computed by a weighted automaton over $S$ but its mirror cannot (for example, the function $w\mapsto\{w\}$ over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ is computed by a weighted automaton over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ but the mirror of this function ($w\mapsto\{\bar{w}\}$) is not).
2.2. Learning with membership and equivalence queries
We are investigating one type of learning done by membership and equivalence queries [4, 8, 7]. In certain circumstances (i.e. by placing conditions on the weighted automaton and the semiring) it is known that one can learn efficiently a weighted automaton computing the function $f$ when provided with finitely many data points $f(w)$ and some steering. Specifically, suppose that there is an oracle who can answer two types of questions about $f$:
  • Membership: when provided with a word $w$: what is the value of $f$ on $w$?
  • Equivalence: when provided with a weighted automaton $\mathcal{H}$: is $f$ the function computed by $\mathcal{H}$? If the oracle answers “no”, it will also give an input word $z$ on which $\mathcal{H}$ performs incorrectly.
A ‘membership’ query is so-named since in the case of the Boolean semiring, this question amounts to asking whether a given word belongs to the rational language recognised by the automaton computing $f$. More generally, a “membership query” (sometimes also called “output query”) provides a mechanism for a learner to gather data points $f(w)$ from which they may attempt to build a hypothesis automaton $\mathcal{H}$ consistent with these data points. An equivalence query provides a mechanism for a learner to test their hypothesis and receive useful feedback, in the form of success or counterexample. Throughout this paper we consider hypothesis automata constructed in a prescribed manner from a partial Hankel matrix, as we shall now explain. What follows has been introduced and studied in the literature; we give proofs of some known results for completeness, putting these into our general framework. Our presentation differs slightly and leads to questions that are answered in Sections 3 and 4 and represent our main contribution.
Hankel matrix:
We will use the notation $\langle x\rangle$ for elements of $S^{\Sigma^{*}}$, and for a word $w\in\Sigma^{*}$ we write $\langle x\rangle_{w}$ to denote the $w$-th component of $\langle x\rangle$. The restriction of $\langle x\rangle$ to a subset $Z\subseteq\Sigma^{*}$ will be denoted by $\langle x\rangle_{Z}$. Given $\langle x\rangle\in S^{\Sigma^{*}}$ and $u\in\Sigma^{*}$, we denote by $\langle xu\rangle$ the element of $S^{\Sigma^{*}}$ defined as the shift of $\langle x\rangle$ by $u$, that is, $\langle xu\rangle_{w}=\langle x\rangle_{uw}$ for all words $w$. Given a function $f:\Sigma^{*}\to S$, the infinite matrix $F$ whose rows and columns are indexed by $\Sigma^{*}$ and with entries $F_{w,w^{\prime}}=f(ww^{\prime})$ is called the Hankel matrix of $f$. If $x$ is a word and $f$ and $F$ are clear from context, we shall write $\langle x\rangle$ to denote the infinite row indexed by $x$ in $F$, viewed as an element of $S^{\Sigma^{*}}$. Note that if $x$ is a word, $f$ a function and $F$ its Hankel matrix, this definition is consistent: $\langle xu\rangle_{w}=F_{xu,w}=f(xuw)=F_{x,uw}=\langle x\rangle_{uw}$. For a subset $X\subseteq S^{\Sigma^{*}}$ and $\langle x\rangle$ in $X$, we will sometimes simplify notation by writing $x\in X$.
Example 2.3.
Consider the function $f$ computed by our running example from Figure 1(a). Its Hankel matrix is given by $F_{\varepsilon,\varepsilon}=-\infty$, $F_{\varepsilon,aw}=|aw|_{a}$, $F_{\varepsilon,bw}=|bw|_{b}$, $F_{aw,w^{\prime}}=|aww^{\prime}|_{a}$ and $F_{bw,w^{\prime}}=|bww^{\prime}|_{b}$ for all words $w,w^{\prime}$.
Closed sets and generating sets:
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) left-closed if for all $\langle x\rangle$ in $X$ and $a\in\Sigma$, $\langle xa\rangle$ is left-generated by $X$; (ii) left-closed on $Z$ if for all $\langle x\rangle$ in $X$ and $a\in\Sigma$, $\langle xa\rangle_{Z}$ is left-generated by the elements of $X$ restricted to $Z$; (iii) row-generating for $f$ if every row of the Hankel matrix of $f$ is left-generated by $X$. By an abuse of notation, we say that a set of words [^1] Using $Q$ as the label for a set of words might seem odd at this point, but the reason behind this will become apparent very shortly. $Q$ is left-closed (resp. left-closed on $Z$, row-generating for $f$) if the set $X=\{\langle q\rangle:q\in Q\}$ has this property.
Observe that for any subset $X\subseteq S^{\Sigma^{*}}$, if $X$ is left-closed and left-generates the row $\langle\varepsilon\rangle$ of the Hankel matrix of $f$ then $X$ is row-generating for $f$ (indeed: since $X$ left generates $\langle\varepsilon\rangle$ we may write $\langle\varepsilon\rangle=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle x\rangle$, with only finitely many coefficients non-zero. Then for all words $u,v\in\Sigma^{*}$ we have $\langle u\rangle_{v}=\langle\varepsilon\rangle_{uv}=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle x\rangle_{uv}$, hence giving $\langle u\rangle_{v}=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle xu\rangle_{v}$. Using the fact that $X$ is left-closed, an easy induction on the length of $u$ then gives that $\langle xu\rangle$ is left-generated by $X$, showing that each row $\langle u\rangle$ is a left-linear combination of elements of $X$.) Conversely, if $X\subseteq S^{\Sigma^{*}}$ is a row-generating set for $f$, then we clearly have in particular that $X$ left-generates the row $\langle\varepsilon\rangle$, and in the case where $X$ is a subset of the rows of the Hankel matrix of $f$, we also have that $X$ is left-closed (since we may view the elements $\langle x\rangle\in X$ in correspondence with words $x\in\Sigma^{*}$, in which case $xa\in\Sigma^{*}$ and hence $\langle xa\rangle$ is left-generated by $X$).
Example 2.4.
Consider again the function $f$ computed by our running example from Figure 1(a) over the semiring $\mathbb{N}_{\rm max}$ whose multiplicative operation $\otimes$ is given by usual addition. The singleton set $\{\langle\varepsilon\rangle\}$ is not left-closed: indeed, $\langle a\rangle$ is not left-generated by this element since $\langle\varepsilon\rangle_{\varepsilon}=-\infty$ but $\langle a\rangle_{\varepsilon}=1$, and hence there is no left-linear combination. The two element set $\{\langle\varepsilon\rangle,\langle a\rangle\}$ is left-closed on $\{\varepsilon\}$, since $\langle b\rangle_{\varepsilon}=\langle ab\rangle_{\varepsilon}=1=0+1=0\otimes\langle a\rangle_{\varepsilon}$ and $\langle aa\rangle_{\varepsilon}=2=1+1=1\otimes\langle a\rangle_{\varepsilon}$. The set $\{\langle\varepsilon\rangle,\langle a\rangle,\langle b\rangle\}$ is left-closed and also row-generating since for $x\in\{a,b\}$, $\langle xw\rangle=|w|_{x}\otimes\langle x\rangle$.
It turns out that if one can find a finite subset $X\subseteq S^{\Sigma^{*}}$ that is left-closed and such that the row $\langle\varepsilon\rangle$ is left-generated by $X$, then one can construct an automaton computing the correct function. We call this a Hankel automaton and describe its construction below. However, in practice, using finitely many membership queries only a finite submatrix of the Hankel matrix can be obtained. A hypothesis automaton is defined in the same way as a Hankel automaton, replacing the finite subset $X$ by a finite set of rows of the Hankel matrix (i.e. indexed by a set of words) restricted to a finite set of columns. For Boolean automata, it is always possible to find a finite set amongst the rows of the Hankel matrix and a finite set of columns such that the corresponding hypothesis automaton computes the target function. The $L^{*}$-algorithm is based on this idea.
Hankel automata and hypothesis automata:
We first set up some notation to describe the construction of hypothesis automaton.
Notation 2.5.
Given a function $f:\Sigma^{*}\rightarrow S$, a finite subset $Q$ of $S^{\Sigma^{*}}$ and a subset $T$ of $\Sigma^{*}$, we define $\Lambda_{Q,T}\subseteq S^{Q}\times S^{Q\times\Sigma\times Q}$ to be the set of tuples $\lambda=(\lambda_{q}^{\varepsilon},\lambda_{q,a,p})_{q\in Q,a\in\Sigma,p\in Q}$ such that:
  • $\langle\varepsilon\rangle_{T}=\bigoplus_{q\in Q}\lambda_{q}^{\varepsilon}\otimes\langle q\rangle_{T}$, and
  • for all $q\in Q$ and all $a\in\Sigma$, $\langle qa\rangle_{T}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle_{T}$.
Remark 2.6.
Note that the set $\Lambda_{Q,T}$ depends on the function $f$ (and hence also the semiring $S$), although we have chosen to suppress this dependence in our notation for readability. An element $\lambda\in\Lambda_{Q,T}$ represents a collection of coefficients witnessing that $\langle\varepsilon\rangle_{T}$ and each of the $\langle qa\rangle_{T}$ (for all $q\in Q$ and all $a\in\Sigma$) is a left-linear combinations of the $\langle q\rangle_{T}$ for $q$ in $Q$. In other words, $\Lambda_{Q,T}$ represents the solution space to the system of $S$-linear equations outlined in Notation 2.5. The set $\Lambda_{Q,T}$ is non-empty if and only if a solution to this system exists; in the language introduced above, if and only if $Q$ is left-closed on $T$ and $\langle\varepsilon\rangle_{T}$ is left-generated by $Q$ restricted to $T$. Note also that if $T\subseteq T’$ ’, then $\Lambda_{Q,T}\supseteq\Lambda_{Q,T’}$. If $T=\Sigma^{*}$, we simply write $\Lambda_{Q}$.
Definition 2.7.
Given a finite subset $Q$ of $S^{\Sigma^{*}}$ and a subset $T$ of $\Sigma^{*}$, for each $\lambda\in\Lambda_{Q,T}$, we define a finite weighted automaton $\mathcal{H}_{Q,T,\lambda}$ with:
  • finite set of states $Q$;
  • initial-state vector $(\lambda^{\varepsilon}_{q})_{q\in Q}$;
  • final-state vector $(\langle q\rangle_{\varepsilon})_{q\in Q}$; and
  • for all $q,p\in Q$ and all $a\in\Sigma$, a transition from $q$ to $p$ labelled by $a$ with weight $\lambda_{q,a,p}$.
If $T=\Sigma^{*}$, we say $\mathcal{H}_{Q,T,\lambda}$ is a Hankel automaton and write simply $\mathcal{H}_{Q,\lambda}$ to reduce notation. If $Q$ is a finite set of rows of the Hankel matrix and $T$ is finite, we say that $\mathcal{H}_{Q,T,\lambda}$ is a hypothesis automaton, and view $Q$ as a finite set of words by identifying with an appropriate index set.
Example 2.8.
Consider the function $f$ over $\mathbb{N}_{\max}$ computed by our running example from Figure 1(a). The set $\{\langle\varepsilon\rangle,\langle a\rangle,\langle b\rangle\}$ clearly left-generates $\langle\varepsilon\rangle$ and, as shown previously, is left-closed. Let $q_{\varepsilon},q_{a}$ and $q_{b}$ denote these rows. Then considering the coefficients $\lambda^{\varepsilon}_{q_{\varepsilon}}=0$, $\lambda_{q_{\varepsilon},a,q_{a}}=\lambda_{q_{\varepsilon},b,q_{b}}=\lambda_{q_{a},b,q_{a}}=\lambda_{q_{b},a,q_{b}}=0$, $\lambda_{q_{a},a,q_{a}}=\lambda_{q_{b},b,q_{b}}=1$ and all non-specified coefficients $\infty$- ∞, the corresponding Hankel automaton would be obtained from the one in Figure 1(a) by replacing the weight of the transitions starting at $q_{\varepsilon}$ by 00, and replacing the final weights by $1$. For $Q=\{q_{\varepsilon},q_{a}\}$ and $T=\{\varepsilon\}$ we have also seen that $Q$ is left-closed on $T$. By taking $\lambda^{\varepsilon}_{q_{\varepsilon}}=0$, $\lambda_{q_{\varepsilon},a,q_{a}}=\lambda_{q_{\varepsilon},b,q_{a}}=\lambda_{q_{a},b,q_{a}}=0$, $\lambda_{q_{a},a,q_{a}}=1$ and all remaining coefficients $\infty$- ∞, one obtains a hypothesis automaton with two states. Notice that this automaton does not compute $f$; it computes $|w|_{a}$ if $w$ starts with an $a$, $|w|_{a}+1$ if $w$ starts with a $b$, and $\infty$- ∞ if $w=\varepsilon$. Note that if $\{\varepsilon,aw\}\subseteq T^{\prime}$ for some word $w$, then $Q=\{q_{\varepsilon},q_{a}\}$ is not left-closed on $T^{\prime}$.
In the case where $S$ is a field the following result can be found in [19]. However, the proof can easily be adapted to the case of arbitrary semirings, and since the details are instructive, we include a short proof.
Theorem 2.9.
Let $f:\Sigma^{*}\rightarrow S$ be a function. There exists a finite subset $Q\subseteq S^{\Sigma^{*}}$ that is left-closed and left-generates $\langle\varepsilon\rangle$ if and only if $f$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$. Specifically, if $Q\subseteq S^{\Sigma^{*}}$ is a finite left-closed set that left-generates $\langle\varepsilon\rangle$, then $\mathcal{H}_{Q,\lambda}$ computes $f$ for all $\lambda\in\Lambda_{Q}\neq\emptyset$, whilst if $f$ is computed by the finite-state weighted automaton $\mathcal{A}$ over $S$, then the set of all $\langle q\rangle$ (where $q$ is a state of $\mathcal{A}$ and $\langle q\rangle_{w}$ is defined as the value of $w$ in $\mathcal{A}$ if $q$ was considered the unique initial state with weight $1_{S}$ and all other weights are the same) is left-closed and left-generates every row of the Hankel matrix of $f$.
Remark 2.10.
Before presenting the proof, we alert the reader to the fact that in Theorem 2.9 the set $Q$ is not assumed to be a set of rows of the Hankel matrix of $f$. In light of Remark 2.6, if $Q$ is a finite set of rows of the Hankel matrix of $f$ and $\Lambda_{Q}$ is non-empty, then for any $\lambda\in\Lambda_{Q}$ the Hankel automaton $\mathcal{H}_{Q,\lambda}$ computes the function $f$. For any finite set $T\subseteq\Sigma^{*}$ we then necessarily have $\lambda\in\Lambda_{Q,T}$, and so (by construction) the hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ constructed from the finite $(Q,T)$ submatrix of the Hankel matrix of $f$, will also compute $f$. The success of Angluin-type algorithms in the existing literature are based on the ability to find (using finitely many membership and equivalence queries): a finite set of words $Q$ such that $\Lambda_{Q}$ is non-empty, together with a finite set of words $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$.
Proof.
Suppose that $Q\subseteq S^{\Sigma^{*}}$ is a finite set that is left-closed and left-generates $\langle\varepsilon\rangle$. By definition, $\Lambda_{Q}$ is non-empty. We prove that if $\lambda$ belongs to $\Lambda_{Q}$ then $\mathcal{H}_{Q,\lambda}$ computes $f$. We first show that for all $q$ in $Q$ and $w$ in $\Sigma^{*}$, $\langle q\rangle_{w}$ is equal to the “weight of $w$ from $q$” that is, the weight computed on input $w$ by the automaton obtained from $\mathcal{H}_{Q,\lambda}$ by imposing the condition that $q$ is the unique initial state with weight $1_{S}$, and keeping all other weights the same. The proof is by induction on the length of $w$. If $w=\varepsilon$, $\langle q\rangle_{\varepsilon}$ is (by construction) equal to the final weight of $q$ in $\mathcal{H}_{Q,\lambda}$ which proves the property. For $a$ in $\Sigma$ and $w$ in $\Sigma^{*}$, $\langle q\rangle_{aw}=\langle qa\rangle_{w}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle_{w}$ since $\lambda\in\Lambda_{Q}$. By induction hypothesis, $\langle p\rangle_{w}$ is the weight of $w$ from $p$, and moreover $\lambda_{q,a,p}$ is (by construction) the weight of the transition from $q$ to $p$ labelled by $a$. So $\langle q\rangle_{aw}$ is indeed the weight of $aw$ from $q$. To conclude the proof, recall that $\langle\varepsilon\rangle$ denotes the row of $\varepsilon$ in the Hankel matrix of $f$, and so by definition of the Hankel matrix of $f$ and the fact that $\lambda\in\Lambda_{Q}$ we have $f(w)=\langle\varepsilon\rangle_{w}=\bigoplus_{q\in Q}\lambda^{\varepsilon}_{q}\otimes\langle q\rangle_{w}$; by our previous observation the latter is then exactly the value computed by $\mathcal{H}_{Q,\lambda}$ on input $w$.
Now suppose that $f:\Sigma^{*}\to S$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$ with set of states $Q$. For each $q\in Q$, let ${}_{q}\mathcal{A}$ (resp. $\mathcal{A}_{q}$) denote the automaton obtained from $\mathcal{A}$ by making $q$ the unique initial (resp. final) state with weight $1_{S}$. Define $\langle q\rangle\in S^{\Sigma^{*}}$ where $\langle q\rangle_{u}$ is equal to the value computed by ${}_{q}\mathcal{A}$ on $u$. Let $R=\{\langle q\rangle\mid q\in Q\}$. First $R$ is left-closed since, by definition, for all words $u\in\Sigma^{*}$, all states $q\in Q$ and all $a\in\Sigma$ we have
$\langle qa\rangle_{u}=\langle q\rangle_{au}=\bigoplus_{p\in Q}(\nu_{p}\otimes\langle p\rangle_{u})$
where $\nu_{p}$ is the weight of the transition labelled by $a$ in $\mathcal{A}$ from $q$ to $p$. Let $w$ be a word in $\Sigma^{*}$ and let $\langle w\rangle$ denote the infinite row of the Hankel matrix of $f$ indexed by $w$. Let $\mu_{q}(w)$ be the value computed by $\mathcal{A}_{q}$ on input $w$. Then, by definition, $\langle w\rangle=\bigoplus_{q\in Q}(\mu_{q}(w)\otimes\langle q\rangle)$, showing that $R$ is row-generating for $f$ and hence in particular $\langle\varepsilon\rangle$ is left-generated by $R$. ∎
Remark 2.11.
In this section we have taken an approach based on “rows”, describing how to construct hypothesis automata utilising finitely many partial rows of the Hankel matrix. Clearly, a dual construction can be made working with the transpose of the Hankel matrix, that is, using partial columns an corresponding notions of right-linear combinations, right semimodules, right-generating sets, right-closed sets, column-generating sets, etc. In the text that follows we shall refer to the analogues of Hankel automata and hypothesis automata constructed by this dual method as co-Hankel automata and co-hypothesis automata. For the readers convenience we outline the details of these constructions and summarise the dual definition to those presented in the main text in Appendix A.
The previous result together with its dual (see Theorem A.3 of the appendix) give the following:
Corollary 2.12.
Let $S$ be a semiring and $f:\Sigma^{*}\rightarrow S$. The following are equivalent:
  1. (1)
    $f$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$;
  2. (2)
    the rows of the Hankel matrix of $f$ lie in a left-subsemimodule of $S^{\Sigma^{*}}$ generated by a finite left-closed set;
  3. (3)
    the columns of the Hankel matrix of $f$ lie in a right-subsemimodule of $S^{\Sigma^{*}}$ generated by a finite right-closed set.
Remark 2.13.
Notice that if $S$ is a field (as in [19]) then the second condition of the previous result implies that the rows of the Hankel matrix must lie in a finitely generated vector space, which in turn forces the vector space spanned by the rows of the Hankel matrix to be finitely generated. Thus if $f$ is a function computed by a finite automaton over a field $S$, one can find finite sets of words $Q,P\subseteq\Sigma^{*}$ such that the rows of the Hankel matrix indexed by $Q$ are row-generating and the columns of the Hankel matrix indexed by $P$ are column-generating. However, over an arbitrary semiring, a subsemimodule of a finitely generated semimodule need not be finitely generated, and so there is no reason to suspect that the left-semimodule generated by the rows of the Hankel matrix is finitely generated. Moreover, it can happen that the left-semimodule generated by the rows of the Hankel matrix is finitely generated, whilst the right-semimodule generated by the columns is not (or vice versa).
Arising questions and structure of the paper:
It is known that for all functions computed by weighted automata over fields there is an efficient way to find $Q$, $T$ and $\lambda$, by membership and equivalence queries such that the hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computes the target function (see for example [8]). For general semirings, the following questions arise:
  • For which functions $f:\Sigma^{*}\rightarrow S$ does there exist a hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computing $f$? We will call the class of all such functions weakly guessable. Clearly, if $f$ is not weakly guessable then any learning algorithm based around making hypotheses of this kind will fail to learn $f$.
  • For the class of weakly guessable functions, is there a learning algorithm using membership and equivalence queries that allows a learner to construct a finite sequence of hypothesis automata, where the last in the sequence computes the target function? In other words: when there exists at least one correct hypothesis, can we devise a strategy to find one? A possible issue here is that even when we have settled upon a suitable set $Q$ with $\Lambda_{Q}\neq\emptyset$, it may be that there is no finite set $T$ such that $\Lambda_{Q,T}\setminus\Lambda_{Q}$ is finite, and there is no obvious general method to discover a “correct” $\lambda$ from the coefficient set in finite time. We call functions $f$ for which there exist finite sets $Q$ and $T$ such that every hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ with $\lambda\in\Lambda_{Q,T}$ computes $f$. There is a naive algorithm (by enumerating all the words for $Q$ and $T$) that permits us to learn the class of guessable functions.
  • For the class of guessable functions, a natural question is whether there is a “fast” algorithm that permits us to learn this class, by making more use of counter-examples, in the style of the successful approaches used for fields and principal ideal domains. These approaches typically alternate between incrementation of the sets $Q$ and $T$, eventually converging on a pair for which $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. A potential problem here is that, in general, one does not know for a given set $Q$ with $\Lambda_{Q}\neq\emptyset$ whether it is worth persevering with incrementation of $T$ or not. We define strongly guessable functions as the ones for which for every suitable set $Q$, there exist a finite set $T$, such that every associated hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ with $\lambda\in\Lambda_{Q,T}$ computes $f$.
In Section 3, we define the classes of weakly guessable, guessable and strongly guessable functions and give several equivalent characterisations for them in Theorem 3.5 and 3.12. We outline a simple learning algorithm for (strongly) guessable functions and discuss further algorithmic issues to be considered including guarantees for termination. In Section 4, we give a general hierarchy for these classes with some specific examples, and give conditions for this hierarchy to collapse. This allows us to have a clearer picture of what happens for several well-studied examples of semirings arising in application areas, where limitations of this approach may be of interest. This opens many questions for further research that are stated in the concluding Section 5.
3. Guessable functions
In this section, we define weakly guessable, guessable, and strongly guessable functions and give characterisations for them.
3.1. Weakly guessable functions
Definition 3.1.
A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly guessable if there exists a finite set $Q$ of rows of the Hankel matrix of $f$ such that $\Lambda_{Q}$ is non-empty.
Observe that by Theorem 2.9, this implies that for all $\lambda\in\Lambda_{Q}$ the automaton $\mathcal{H}_{Q,\lambda}$ computes $f$ and hence for any non-empty $T$, there exists a hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computing $f$ since $\Lambda_{Q}\subseteq\Lambda_{Q,T}$. So this definition means that there is “some hope” that an algorithm à la Angluin can find a correct automaton, but does not really give a handle on what these functions look like. The following definition turns out to provide a syntactic charaterisation of these functions in terms of automata and links to residual automata. A literal automaton is a residual automaton, as defined for example for non-deterministic or nominal automata [20, 28] with extra conditions to deal with the weights. It is very close to the reduced residual automata in the stochastic settings defined in [15], with a slight difference in the way the weights are handled - the reason being that, while non-zero probabilities have an inverse for the multiplication, this is not systematically the case in general semirings. We use the terminology literal automata to avoid any confusion with [15], our choice of terminology reflecting the fact that each state of a literal automaton can be labelled by a word with the property that any run from the initial state labelled by that word terminates in the given state.
Definition 3.2.
A weighted automaton over $S$ with set of states $Q$ is said to be literal if it has a unique initial state with weight $1_{S}$ and there exists a prefix-closed set of words $W\subseteq\Sigma^{*}$ and a bijection $\sigma:Q\rightarrow W$ such that for all $q\in Q$:
  • there is a unique run, $\gamma_{q}$, starting at the initial state and labelled by $\sigma(q)$,
  • each transition in $\gamma_{q}$ has weight $1_{S}$,
  • $\gamma_{q}$ terminates at state $q$.
This means that in a literal automaton $\mathcal{A}$, each state $q$ can be characterised by a word $\sigma(q)$ that leads uniquely to that state (and with weight $1_{S}$), and hence the set $\{\langle q\rangle\mid q\text{ is a state of }\mathcal{A}\}$ defined in Theorem 2.9 is exactly the set of rows of the Hankel matrix of $\mathcal{A}$ indexed by the words $\sigma(q)$.
Finally, we link the functions computed by literal automata to a weaker variant of a property given in [21], demonstrating that this corresponds exactly to the weakly guessable functions.
Definition 3.3.
Given a function $f:\Sigma^{*}\to S$ and $F$ its Hankel matrix, $f$ satisfies the weak ascending chain condition if for all chains of left-semimodules $X_{0}\subseteq X_{1}\subseteq X_{2}\subseteq\ldots$ such that
  • each $X_{i}$ is generated by a finite set of rows of $F$;
  • each row of $F$ belongs to some $X_{i}$
there exists $n$ such that for all $m\geq n$, $X_{m}=X_{n}$.
Remark 3.4.
The terminology “weak” in the previous definition reflects the fact that we only consider certain chains of left-semimodules. A left-semimodule is said to be Noetherian if all chains of left-subsemimodules satisfy the ascending chain condition. Thus, if every left-semimodule generated by a finite set of rows of $F$ is Noetherian, then $f$ satisfies the weak ascending chain condition above, but not conversely (in general, a submodule of a finitely generated semimodule need not be finitely generated, nor need a subsemimodule of a semimodule generated by rows of the Hankel matrix also be generated by rows of the Hankel matrix).
Theorem 3.5.
Given a function $f:\Sigma^{*}\to S$, the following assertions are equivalent:
  1. (1)
    $f$ is weakly guessable;
  2. (2)
    there exists a literal automaton computing $f$;
  3. (3)
    there exists a finite set of rows of the Hankel matrix of $f$ that is row-generating for $f$;
  4. (4)
    $f$ satisfies the weak ascending chain condition.
Remark 3.6.
Comparing Theorem 3.5 (characterising weakly guessable functions) with Theorem 2.9 (characterising functions computed by a finite weighted automata), notice that point (3) of Theorem 3.5 states that there is a finite subset of the rows of the Hankel matrix that left-generates every row of the Hankel matrix, whilst Theorem 2.9, states that the rows of the Hankel matrix can be generated by finitely many elements of $S^{\Sigma^{*}}$ (which need not themselves be rows of the Hankel matrix; see also Remark 2.10).
Proof.
(1. implies 2.) Since $f$ is weakly guessable there is a finite set of rows $Q$ of the Hankel matrix such that $\Lambda_{Q}$ is non-empty; by a slight abuse of notation, we shall view $Q$ as a subset of $\Sigma^{*}$, identifying a row of a Hankel matrix with the word labelling that row. Fix $\lambda\in\Lambda_{Q}$ and recall that (by Theorem  2.9) $\mathcal{H}_{Q,\lambda}$ computes $f$. Let $Q^{\prime}$ denote the set of all prefixes of words in $Q$, and define $\mu=(\mu^{\varepsilon}_{q},\mu_{q,a,p})_{p,q\in Q^{\prime},a\in\Sigma}$ as follows:
  • $\mu^{\varepsilon}_{q}=1_{S}$ if and only if $q=\varepsilon$ and $0_{S}$ otherwise,
  • if $q\in Q^{\prime}$, $a\in\Sigma$ and $qa\in Q^{\prime}$, then $\mu_{q,a,qa}=1_{S}$, and $\mu_{q,a,p}=0_{S}$ for all other $p\in Q^{\prime}$;
  • if $q\in Q$, $a\in\Sigma$ and $qa$ is not in $Q^{\prime}$, then $\mu_{q,a,p}=\lambda_{q,a,p}$ for all $p\in Q$ and $\mu_{q,a,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$;
  • if $q$ in $Q^{\prime}\setminus Q$, $a$ in $\Sigma$ and $qa$ is not in $Q^{\prime}$, then for each $p\in Q^{\prime}$ we define $\mu_{q,a,p}$ by induction on the length of $q$, as follows:
    If $q=\varepsilon$, then for all words $w$ we have $\langle\varepsilon\rangle_{w}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle r\rangle_{w}$ and
    $\langle a\rangle_{w}=\langle\varepsilon\rangle_{aw}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle r\rangle_{aw}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle ra\rangle_{w}=\bigoplus_{r,p\in Q}\lambda^{\varepsilon}_{r}\otimes\lambda_{r,a,p}\otimes\langle p\rangle_{w}.$
    In this case we set: $\mu_{\varepsilon,a,p}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\lambda_{r,a,p}$ for all $p\in Q$ and $\mu_{\varepsilon,a,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$. If $q=ra$ and $\mu_{r,a,p}$ has already been defined for all $p\in Q^{\prime}$ in such a way that expresses $\langle ra\rangle$ as a left-linear combination of the rows $Q$ (i.e. such that $\mu(r,a,p)=0_{S}$ if $p$ is not in $Q$), then for all words $w$ and all $b\in\Sigma$ we have
    $\langle qb\rangle_{w}$ $=$ $\langle rab\rangle_{w}=\langle ra\rangle_{bw}=\bigoplus_{p\in Q}\mu_{r,a,p}\otimes\langle p\rangle_{bw}$
    $=$ $\bigoplus_{p\in Q}\mu_{r,a,p}\otimes\langle pb\rangle_{w}=\bigoplus_{p,s\in Q}\mu_{r,a,p}\otimes\lambda_{p,b,s}\otimes\langle s\rangle_{w}.$
    In this case we set $\mu_{ra,b,p}=\bigoplus_{s\in Q}\mu_{r,a,p}\otimes\lambda_{p,b,s}$ for all $p\in Q$ and $\mu_{ra,b,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$.
By construction we then have that $\mu$ belongs to $\Lambda_{Q^{\prime}}$ and hence by Lemma 2.9, $\mathcal{H}_{Q^{\prime},\mu}$ computes $f$. The automaton $\mathcal{H}_{Q^{\prime},\mu}$ is also literal by construction, which concludes the proof.
(2. implies 3.) Let $\mathcal{A}$ be a literal automaton computing $f$ and $Q$ its set of states. By Theorem 2.9, the set $R=\{\langle q\rangle\mid q\in Q\}$, where $\langle q\rangle_{w}$ is defined as the weight of $w$ from $q$, is row-generating for $f$. Moreover, by definition of literal automaton, there is a finite set of words $\{w_{q}\mid q\in Q\}$ such that for all $q$, $\langle q\rangle=\langle w_{q}\rangle$.
(3. implies 4.) Let $W\subseteq\Sigma^{*}$ such that $\{\langle w\rangle\mid w\in W\}$ is row-generating and let a sequence of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ as in the definition of the weak ascending chain condition. Then there is $n$ such that for all $m\geq n$, all the rows $\langle w\rangle$ for $w$ in $W$ are in $X_{m}$, and since they are row-generating, $X_{m}=X_{n}$.
(4. implies 1.) For all positive integers $i$, let $X_{i}$ be the left-semimodule generated by the rows indexed by words of length at most $i$. Then the $X_{i}$ satisfy the hypothesis of the weak acending chain condition and hence there is $n$ such that for all $m$, $X_{m}=X_{n}$. Let $Q$ be the set of words of length at most $n$. Necessarily the set of rows indexed by $Q$ is row-generating (and hence left-closed, and left-generates $\langle\varepsilon\rangle$) so there exists some $\lambda\in\Lambda_{Q}$, which concludes the proof. ∎
Remark 3.7.
As noted in Remark 2.11, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions which we refer to as weakly co-guessable, column generating sets, and weak co-ascending chain condition: see Appendix A for full details. We say that an automaton is co-literal if its mirror is literal. With these definitions in place, one can prove the following entirely analogous result.
Theorem 3.8.
Given a function $f:\Sigma^{*}\to S$, the following assertions are equivalent:
  1. (1)
    $f$ is weakly co-guessable;
  2. (2)
    there exists a co-literal automaton computing $f$;
  3. (3)
    there exists a finite set of columns of the Hankel matrix of $f$ that is column-generating for $f$;
  4. (4)
    $f$ satisfies the weak co-ascending chain condition.
Remark 3.9.
The semirings where all functions computed by finite state weighted automata satisfy the weak ascending and co-ascending chain conditions are therefore the semirings for which all functions computed by finite state weighted automata are both weakly guessable and weakly co-guessable (by Theorem 3.5 and its dual Theorem 3.8). This means that over such a semiring all functions computed by a finite weighted automaton have the potential to be learnt, provided one can make a lucky guess for $Q$, $T$ and $\lambda$. This is in particular the case for $\mathbb{B}$ (or indeed any finite semiring), $\mathbb{R}$ (or indeed, any field), and $\mathbb{Z}$ (or indeed, any principal ideal domain c.f.  [21]). More generally, if $R$ is any left-Noetherian ring, then all finitely generated left $R$-modules are Noetherian, and hence by Remark 3.4, all functions computed by finite state weighted automata over $R$ are weakly guessable. This covers the case of finite rings, fields, principal ideal domains, and division rings.
3.2. Guessable and strongly guessable functions
The definition of weakly guessable functions ensures the existence of some finite sets $Q,T$ and coefficients $\lambda\in\Lambda_{Q,T}$ that together determine a correct hypothesis automaton. In other words, functions that are not weakly guessable are those for which all hypothesis automata are incorrect, and so (in fairness to the learner!) one should immediately exclude all such functions from consideration if one seeks to learn by membership and equivalence in the style of Angluin. For weakly guessable functions, the problem then becomes: how to find a suitable $Q,T$ and $\lambda$. Clearly $Q$ must be a row-generating set for the full Hankel matrix; but is it always possible to find such a set from finitely many membership and equivalence queries? Also, even if a suitable row generating set $Q$ is found, there is no guarantee that accompanying $T$ and $\lambda$ can be found in finite time. Guessable functions are those for which there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. Strongly guessable functions are those for which for every finite row-generating set $Q$ there exists a finite subset $T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. In the following definition, the reader should recall from Remark 2.6 that the sets $\Lambda_{Q,T}$ depend upon the function $f$ (and hence also the semiring $S$).
Definition 3.10.
A function $f:\Sigma^{*}\rightarrow S$ is guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is guessable. Furthermore, we say that a guessable function $f$ is strongly guessable if for all finite sets $Q\subseteq\Sigma^{*}$ with $\Lambda_{Q}\neq\emptyset$ there exists a finite set $T\subseteq\Sigma^{*}$ such that $(Q,T)$ witnesses that $f$ is guessable.
It is clear from the definitions that every strongly guessable function is guessable, and every guessable function is weakly guessable. Moreover, observe that, by Theorem 2.9, if the pair $(Q,T)$ witnesses that $f$ is guessable, then all hypothesis automata over $Q$ and $T$ compute $f$. We now give a weaker variant of another property given in [21] to characterise guessable functions.
Definition 3.11.
Let $Q$ be a finite subset of $\Sigma^{*}$. A total $Q$-chain is an infinite descending chain of solution spaces $\Lambda_{Q,T_{0}}\supsetneq\Lambda_{Q,T_{1}}\supsetneq\Lambda_{Q,T_{2}}\supsetneq\cdots$ with $T_{i}\subseteq T_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $T_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is row-bound if there exists a finite set of words $Q$ such that $\Lambda_{Q}\neq\emptyset$ and there is no total $Q-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly row-bound if for every finite set of words $Q$ such that $\Lambda_{Q}\neq\emptyset$ there is no total $Q$-chain.
The above definition is to be linked with the progress measure defined in [21, Definition 11]. The existence of a progress measure for a Hankel matrix implies that the corresponding function is strongly row-bound. However, the converse might not be true. A progress measure implies a uniform bound, for all $Q$, on the number of elements in a total $Q$-chain, but for a (strongly) row-bound function, even if no such chain can be infinite, its size might depend on the set of words $Q$. The progress measure is a stronger assumption, and one of the sufficient condition for termination of a learning algorithm [21], while the property of being (strongly) row-bounded pinpoints more precisely a characterisation of (strongly) guessable functions. We give a variant of the notion of progress measure in the next subsection.
Proposition 3.12.
Let $f:\Sigma^{*}\to S$ be a weakly guessable function. Then $f$ is (strongly) guessable if and only if $f$ is (strongly) row-bound.
Proof.
Suppose that $Q,T$ are finite sets of words such that $\Lambda_{Q,T}=\Lambda_{Q}$. Let $(T_{i})_{i\geq 0}$ be a sequence of subsets of $\Sigma^{*}$ with $T_{i}\subseteq T_{i+1}$ for all $i$ and such that $\bigcup_{i}T_{i}=\Sigma^{*}$. By definition, there exists some set $T_{j}$ such that $T\subseteq T_{j}$. Then for all $i\geq j$ we have $T\subseteq T_{j}\subseteq T_{i}$. Since $\Lambda_{Q,T}=\Lambda_{Q}$ it then follows that $\Lambda_{Q,T_{i}}=\Lambda_{Q,T_{j}}=\Lambda_{Q,T}$ for all $i\geq j$. Thus there can be no total $Q$-chains. It now follows that if $f$ is guessable then $f$ is row-bound, whilst if $f$ is strongly guessable, then it must be strongly row-bound.
Suppose now that $f$ is (strongly) row-bound. Let $Q$ be a finite set of words such that $\Lambda_{Q}\neq\emptyset$ and there is no total $Q$-chain. Then taking $T_{i}$ to be the finite set of words of length at most $i$ for each positive integer $i$, we have that $T_{i}\subseteq T_{i+1}$ and $\bigcup_{i}T_{i}=\Sigma^{*}$ and so there must exist $n$ such that for all $m\geq n$, $\Lambda_{Q,T_{m}}=\Lambda_{Q,T_{n}}$. Since the $T_{i}$ enumerate all words, we must therefore also have $\Lambda_{Q,T_{n}}=\Lambda_{Q}$. Hence for any $\lambda$ in $\Lambda_{Q,T_{n}}$, the hypothesis automaton over $Q,T_{n}$ and $\lambda$ computes $f$. Thus, if $f$ is (strongly) row-bound, we conclude that $f$ is (strongly) guessable. ∎
Remark 3.13.
As in Remarks 2.11 and 3.7, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions of co-guessable, strongly co-guessable: see Appendix A for full details. With these definitions, one can prove the following entirely analogous result.
Proposition 3.14.
Let $f:\Sigma^{*}\to S$ be a weakly co-guessable function. Then $f$ is (strongly) co-guessable if and only if $f$ is (strongly) column-bound.
Remark 3.15.
Let $S$ be a field and $f:\Sigma^{*}\rightarrow S$ a function computed by a finite automaton with weights from $S$. By Remark 2.13 there exist finite sets of words $Q,P\subseteq\Sigma^{*}$ such that $Q$ is row-generating for $f$ and $P$ is column-generating for $f$. It is straightforward to check that for any such sets $Q$ and $P$ we have $\Lambda_{Q,P}=\Lambda_{Q}\neq\emptyset$ (see Lemma 4.1 below for a generalisation of this fact) and hence $f$ is strongly guessable (and by a dual argument, $f$ is strongly co-guessable).
3.3. Learning algorithms
As mentioned earlier, if one knows that the target function is weakly guessable, that is to say that there is a hypothesis automaton that computes it, the question becomes to find some suitable $Q,T$ and $\lambda$. If for each pair $Q,T$ the set $\Lambda_{Q,T}$ is finite (as is the case for any finite semiring), then by enumerating the set of pairs $(Q,T)$, one can construct a naive learning algorithm for weakly guessable functions given in Figure LABEL:fig:algoweak; we say that such functions have finite solvability.
Similarly, by enumerating the set of pairs $(Q,T)$, one can construct (regardless of the cardinalities of the sets $\Lambda_{Q,T}$) the naive learning algorithm for guessable functions given in Figure LABEL:fig:algoguess.
For strongly guessable functions, Figure LABEL:fig:algostrong gives an alternative algorithm in which $Q$ and $T$ are enumerated and incremented alternately.
The applicability of each of these algorithms is dependent on decidability of equivalence queries (although, in practice, it may be possible to approximate equivalence queries using many membership queries) and solvability of general finite $S$-linear systems of equations. Indeed, a requirement for the algorithm to function is to be able to determine whether left-linear combinations exist, and when this is the case, to be able to find at least one such. This is given in the solvability condition.
Definition 3.16.
A function $f$ is solvable if there is an algorithm that, for any finite sets $Q$, $T$, can determine whether $\Lambda_{Q,T}\neq\emptyset$ and compute at least one element of $\Lambda_{Q,T}$ in that case.
Clearly, if $S$ is a semiring where there exists an algorithm which can take as input an arbitrary system of linear equations over $S$ and compute a solution, then all functions over the semiring are solvable. We note that in [21] such semirings were called solvable. It is clear that all finite semirings are solvable (a very naive algorithm would be to try all possibilities) and fields where the basic operations are computable are also solvable (via Gaussian elimination). In [21] the authors provide a sufficient condition for a principal ideal domain to be solvable. Turning to other examples, we note for example that over $\mathbb{Z}_{\rm max}$ there is a straightforward (and well-known) method to determine whether a system of the form $A\otimes x=b$ has a solution. Indeed, if $x$ is to be a solution, we must in particular have that $A_{i,j}+x_{j}\leq b_{i}$ for all $i,j$. If $A_{i,j}=-\infty$, then this inequality holds automatically, otherwise, we require that $x_{j}\leq b_{i}-A_{i,j}$. In order to solve the inequality, for each $j$ such that there exists at least one $i$ with $A_{i,j}\neq-\infty$ we could set $x_{j}={\rm min}\{b_{i}-A_{i,j}:A_{i,j}\neq-\infty\}$, and set all remaining $x_{j}$ equal to 00. It is clear that this is a computable (we use finitely many operations of subtraction and comparison on finitely many entries drawn from $\mathbb{Z}_{\rm max}$) solution to the inequality $A\otimes x\leq b$. In some sense it can be viewed as a “greatest” solution since the entries $x_{j}$ where there exists $A_{i,j}\neq-\infty$ cannot be taken any larger (of course, if there are any columns of $A$ where the entries are all equal to $\infty$- ∞, the corresponding entry of $x$ is irrelevant and our choice of taking $x_{j}=0$ was completely arbitrary). Furthermore, it can also be seen that either the solution to the inequality is a solution to the equation $A\otimes x=b$, or else no solution to the equation $A\otimes x=b$ exists. (Indeed, if the solution to the inequality we constructed does not satisfy the equation, then we must have that for some $j$ the strict inequality holds: ${\rm max}(A_{i,j}+x_{j})<b_{i}$, meaning that either all $A_{i,j}$ are equal to $\infty$- ∞ and so there is no hope to satisfy the equation ${\rm max}(A_{i,j}+x_{j})<b_{i}$, or else in order to attain $A_{i,j}+x_{j}=b_{i}$ we require that one of the $x_{j}$ with $A_{i,j}$ must take a larger value, which as we have seen, is not possible.) Thus $\mathbb{Z}_{\rm max}$ (and likewise $\mathbb{N}_{\rm max}$) is solvable.
We also note that each of the naive algorithms outlined above do not make any use of the data provided by the counterexamples. We do not develop further algorithms here: as mentioned in the introduction, our main aim in this paper is to picture the landscape for active learning in arbitrary semirings, and demonstrate the limitations. Our intention is to lay the foundations for future progress, indicating clear avenues for future investigation. However, to conclude this section we reproduce, in Figure LABEL:fig:algo, the general algorithm given in [21], and discuss the relationship to the present work.
In [21], an example showing that this algorithm does not always terminate is given. This example corresponds to the automaton given in Figure 1(d), that (as we shall see in the next section) is not weakly guessable. Since the algorithm can only guess hypothesis automata, it is clear that the algorithm will not terminate on any function that is not weakly guessable. Moreover, if the function is not guessable, then (by definition) for each finite set of words $Q$ with $\Lambda_{Q}\neq\emptyset$ one has that $\Lambda_{Q}\neq\Lambda_{Q,T}$ for all finite sets of words $T$. This in turn means that there is no guarantee that the algorithm will terminate, since it could be that in every test step the chosen $\lambda$ generates an incorrect hypothesis automaton.
In [21], three conditions are requested to ensure termination and correctness of the algorithm in Figure LABEL:fig:algo: the ascending chain condition, the progress measure and the solvability condition. Below we give slight variants on these conditions, that ensure that the proof given in  [21] still holds (and any function satisfying these conditions will be guessable). The following strong ascending chain condition ensures that the closure step always terminates in finite time for a fixed $T$ and that it cannot be taken infinitely many times for different $T$’s.
Definition 3.17.
A function $f:\Sigma^{*}\to S$ with Hankel matrix $F$ satisfies the strong ascending chain condition if for all sequences of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ generated by finite sets of rows of $F$, such that $X_{0}\subseteq X_{1}\subseteq X_{2}\subseteq\cdots$ there is $n$ such that for all $m\geq n$, $X_{m}=X_{n}$.
Meanwhile the following progress measure ensures that one cannot go through the Test step infinitely many times without changing the set $Q$ in the closure step.
Definition 3.18.
Let $f$ be a function computed by a weighted automaton over $S$, let :=$\mathcal{L}:=\{(Q,T):Q,T\mbox{ are finite sets of words and $Q$ is left-closedon $T$}\}$. We say that $f$ has a progress measure if there is a totally ordered set $(\mathbb{K},\succ)$ with no infinite decreasing chain and a function $\mu:\mathcal{L}\rightarrow\mathbb{K}$ with the property that whenever $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$ we have $\mu(Q,T)\succ\mu(Q,T^{\prime})$.
It is clear from the definitions that if $f$ has a progress measure then it must be strongly row-bound. Moreover, with these definitions, the following proposition now derives in a similar manner to results presented in [21].
Proposition 3.19.
[21] If a function $f$ is solvable, satisfies the strong ascending chain condition and has a progress measure, then the learning algorithm from Figure LABEL:fig:algo terminates and returns an automaton computing $f$.
4. Hierarchy of learnable functions
Figure 6. General hierarchy.
In the previous section, we have defined the classes of strongly guessable, guessable, and weakly guessable functions, and in Appendix A the dual notions of strongly co-guessable, co-guessable, and weakly co-guessable functions. We obtain a general hierarchy depicted in Figure 6. Next we show that parts of this hierarchy collapse in some natural settings, such as for commutative semirings (see Figure 7 below) and give examples to demonstrate that certain of these containments can be strict. The examples are drawn from various well-studied semirings discussed in Section 2, but one could obtain a unified semiring separating these classes by considering their product with component-wise operations.
4.1. Commutative semirings
For commutative semirings we have the following result, giving the hierarchy depicted in Figure 7.
Figure 7. Hierarchy for commutative semirings
Lemma 4.1.
Let $S$ be a commutative semiring and let $f:\Sigma^{*}\rightarrow S$ be computed by a weighted automaton over $S$. If $f$ is weakly guessable and weakly co-guessable, then $f$ is strongly guessable and strongly co-guessable.
Proof.
Suppose that $f$ is weakly guessable and weakly co-guessable and let $F$ be the Hankel matrix of $f$. We first show that $f$ is strongly guessable. Since $f$ is weakly co-guessable, we may fix a finite set of words $T$ such that the columns indexed by $T$ are column-generating. We will prove that for any finite set of words $Q$ with $\Lambda_{Q}\neq\emptyset$ we have $\Lambda_{Q}=\Lambda_{Q,T}$.
Thus let $Q$ be arbitrary with the property that $\Lambda_{Q}\neq\emptyset$ and let $\lambda\in\Lambda_{Q,T}$. Then for all $q,p\in Q$, all $a\in\Sigma$ and all $t\in T$ we have: $F_{qa,t}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,t}$. By our assumption on $T$ we know that for all words $w$ and all $t\in T$ there exist $\mu_{w,t}\in S$ such that $F_{u,w}=\bigoplus_{t\in T}\mu_{w,t}\otimes F_{u,t}$ holds for all words $u$. In particular, for all $w\in\Sigma^{*}$, we have:
$F_{qa,w}$ $=$ $\bigoplus_{t\in T}\mu_{w,t}\otimes F_{qa,t}=\bigoplus_{t\in T}\mu_{w,t}\otimes(\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,t})$
$=$ $\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes(\bigoplus_{t\in T}\mu_{w,t}\otimes F_{p,t})$
$=$ $\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,w}.$
In other words, $\langle qa\rangle=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle$, giving $\lambda\in\Lambda_{Q}$, as required. This demonstrates that $f$ is strongly guessable.
Since $f$ is weakly guessable and co-guessable, then $\bar{f}$ is also weakly guessable and weakly co-guessable. The above argument therefore yields that $\bar{f}$ is strongly guessable and so $f$ is strongly co-guessable. ∎
We conclude this section with the following useful observations concerning commutative semirings:
Lemma 4.2.
If $S$ is a commutative semiring then a function is weakly guessable (respectively, guessable, strongly guessable) if and only if its mirror is weakly co-guessable (respectively, co-guessable, strongly co-guessable).
We next aim to show that for commutative semirings almost all inclusions depicted in Figure 7 can be strict.
4.2. Examples for the strict hierarchy.
The full proofs of the membership of the following functions to certain parts of the hierarchy are given in Appendix B. We give here a quick overview.
For some semirings $S$, there are functions $f:\Sigma^{*}\rightarrow S$ that are neither weakly guessable, nor weakly co-guessable. We show that this is the case for $f_{1}$ computed by the automaton given in Figure 1(b), whether viewed over $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$, $f^{\prime}_{1}$ computed by the automaton given in Figure 1(d), whether viewed over $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$, and $f^{\prime\prime}_{1}$ computed by the automaton in Figure 1(f). There are also functions that are weakly guessable but neither guessable nor weakly co-guessable. We show that this is the case for $f_{2}$ computed by the automaton given in Figure 1(a), whether viewed as an automaton over $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$. It then follows that the mirror function $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable.). Finally, we show that the functions $f_{3}$ computed by the automaton depicted in Figure 1(a) (over $\mathbb{N}_{\max}$), $f^{\prime}_{3}$ computed by the automaton from Figure 1(c) and $f^{\prime\prime}_{3}$ computed by the automaton from Figure 1(e) are all strongly guessable but not weakly co-guessable. The functions $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable. It is in general easy to find examples at the intersection of all the classes: for commutative semirings, any function computed by an automaton that is both literal and co-literal lies in this intersection. A trivial example of such function would be the constant function $f_{4}$ mapping all words to $1_{S}$.
4.3. Collapse in some familiar semirings
In many well-studied semirings, we note that the hierarchy collapses further, due to the fact that certain ascending and descending chain conditions are satisfied. It is clear from the results of previous section that the following two properties characterise (A) when all functions computed by a finite state weighted automaton over $S$ are weakly guessable and weakly co-guessable (as in Figure 8(a)) and (B) when all weakly guessable functions are strongly guessable and all weakly co-guessable functions are strongly co-guessable (as in Figure 8(b)):
  • Property A: all functions computed by finite state weighted automata over $S$ satisfy the weak ascending and weak co-ascending chain condition;
  • Property B: all functions computed by finite state weighted automata over $S$ which satisfy the weak ascending chain condition are strongly row-bound and all functions computed by finite state weighted automata over $S$ which satisfy the weak co-ascending chain condition are strongly column-bound.
(a) Property A
(b) Property B
Figure 8. Hierarchies for semirings satisfying properties A and B.
Clearly if both the conditions depicted in Figures 8(a) and 8(b) hold, we have that all functions computed by finite weighted automata are strongly guessable. Note that this is the case in the semirings $\mathbb{Z}$, $\mathbb{R}$ and $\mathbb{B}$ (as special cases of principal ideal domains, fields and finite semiring respectively) for example. Further, the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$ and $\mathcal{P}_{\rm fin}(\Sigma^{*})$ satisfy Property B:
Lemma 4.3.
Every weakly guessable (respectively, weakly co-guessable) function over $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$ is strongly row-bound (respectively, column-bound).
Proof.
We give the proof for rows; he proof for columns is dual. Let $S$ be one of the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$. Let $f$ be a weakly guessable function computed over $S$, and $F$ its Hankel matrix. Let $Q$ be such that $\Lambda_{Q}$ is non-empty. First, note that, for some fixed $T$, if no row indexed by $Q$ and restricted to $T$ is fully $0_{S}$, then in the three cases, $\Lambda_{Q,T}$ is finite. Let $Z_{Q,T}=\{q\in Q\mid\langle q\rangle_{T}\mbox{ is a }0_{S}\mbox{ row}\}$, $Q^{\prime}=Q\setminus Z_{Q,T}$ and $N_{Q^{\prime},T}$ the set of tuples $(\lambda_{p}^{\varepsilon},\lambda_{q,a,p})_{q\in Q,a\in\Sigma,p\in Q^{\prime}}$ from $\Lambda_{Q,T}$ where $p$ is restricted to $Q^{\prime}$. Clearly $Z_{Q,T}$ is finite and its size is bounded above by the one of $Q$. Notice that $N_{Q^{\prime},T}$ is also a finite set, since each $\lambda_{q,a,p}$ where $p\in Q^{\prime}$ can only take a finite number of different values. Suppose that $T\subsetneq T^{\prime}$ with $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$. It is clear from the definition that $Z_{Q,T}\supseteq Z_{Q,T^{\prime}}$, and either this containment is strict, or if not, then $N_{Q^{\prime},T}\supsetneq N_{Q^{\prime},T^{\prime}}$. Since both sets have a finite size, there cannot be an infinite descending chain as in the definition. ∎
We remark that in semirings where the concepts of weakly guessable and strongly guessable (resp. weakly co-guessable are strongly co-guessable) coincide (as is the case for fields and finite semirings, for example), this is inherited by subsemirings:
Lemma 4.4.
Let $S^{\prime}$ be a subsemiring of $S$. If every weakly guessable function over $S$ is strongly guessable, then every weakly guessable function over $S^{\prime}$ is also strongly guessable. In particular, if $S$ is a subsemiring of a field (e.g. an integral domain), then every weakly guessable function over $S$ is strongly guessable.
Proof.
Suppose that $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable. Thus there is a literal automaton $\mathcal{A}$ with weights in $S^{\prime}$ that computes $f$. Viewed as a function over $S$, we also have that $f$ is weakly guessable (since $\mathcal{A}$ is a literal automaton with weights in $S$ computing $f$) and hence by assumption, strongly guessable.
Suppose for contradiction that $f$ not strongly guessable over $S^{\prime}$. Then there exists a finite set $Q$ such that $\Lambda_{Q}\neq\emptyset$ and for all finite sets $T$ there exists $\lambda_{T}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$. However, noting that the coefficients $\lambda_{T}$ lie in $S$, this directly contradicts that $f$ is strongly guessable over $S$.
The final statement follows from Remark 3.9 together with the fact that any integral domain embeds into a field (its field of fractions). ∎
The table in Figure 9 demonstrates that there are interesting and important semirings (including the natural numbers) where the properties of weakly guessable and strongly guessable coincide (i.e. Property B holds) whilst there exist functions that are not weakly guessable (i.e. Property A does not hold). Since the property of being weakly guessable is characterised in terms of automata by Theorem 3.5, over such semirings it is possible for the ‘teacher’ to select a guessable function to begin with. Finally, we say that semirings in which all six classes of guessability collapse have property $D$. Clearly if Properties A and B both hold, then D must also hold. Conversely, if Property B does not hold, then Property D cannot hold. We do not have an example where Propoerty D holds but Property A does not.
A B D
$\mathbb{B}$ Yes Yes Yes
$\mathbb{R}$ Yes Yes Yes
$\mathbb{Z}$ Yes Yes Yes
$\mathbb{N}$ No ($f^{\prime}_{1}$) Yes No ($f^{\prime}_{3}$)
$\mathbb{R}_{\geq 0}$ No ($f^{\prime}_{1}$) Yes No ($f_{5}$)
$\mathcal{P}_{\rm fin}(\Sigma^{*})$ No ($f^{\prime\prime}_{1}$) Yes No ($f^{\prime\prime}_{3}$)
$\mathbb{N}_{\max}$ No ($f_{1}$) Yes No ($f_{3}$)
$\mathbb{Z}_{\max}$ No ($f_{1}$) No ($f_{2}$) No
$\mathbb{R}_{\max}$ No ($f_{1}$) No ($f_{2}$) No
Figure 9. Summary of collapse of classes in familiar semirings.
5. Conclusion and open questions
We have pictured a landscape for learning weighted automata on any semiring and shown that an approach à la Angluin can only work in general on a restricted class of functions. This paper opens many questions, both general and specific. First, for functions that are not weakly guessable, one can ask whether there is still a way to find appropriate states (not from rows of the Hankel matrix) by strengthening the queries. Next, for weakly guessable functions, one can ask whether there is a way to design strategies to update $Q$ and $T$ (other than closure and adding suffixes of counter-examples to $T$) that would ensure to find a suitable hypothesis automaton. In particular, is there a way to choose “cleverly” the coefficients of the linear combinations under consideration? For example, the algorithm of [21] can fail for functions that are not guessable (since there is no condition on how the coefficients are chosen). For guessable functions, a naive way to ensure termination is to make sure to enumerate all the words in $T$. But is there a better way? And if so, it would be interesting to see if the class of functions learnable by a polynomial time algorithm can be characterised. Regarding the general hierarchy of functions, there are several classes for which we neither have an example of non-emptiness or an argument for collapse. It would also be of interest to understand how termination of the algorithm from [21] relates to strong guessability.
Acknowledgment
This work has been supported by the EPSRC grant EP/T018313/1 and a Turing-Manchester Exchange Fellowship, The Alan Turing Institute. We also thank Nathanaël Fijalkow for useful discussions on the topic.
Appendix A Dual approach using columns
In this section, we present (for clarity) the natural dual definitions to those provided in Section 2 and Section 3, working with the transpose of the Hankel matrix. Thus, we work with (partial) columns instead of (partial) rows and consider right multiplication instead of left multiplication. As a result, the role of final and initial states is swapped in our constructions. $S$ still denotes a semiring and $\Sigma$ an alphabet.
Right semimodules, closed sets and right generating sets:
Given a set $Y$, an element $x$ of $S^{Y}$ and $\lambda$ in $S$, $x\otimes\lambda$ denotes the element of $S^{Y}$ computed from $x$ by multiplying every component of $x$ by $\lambda$ on the right. Given a set $X=\{x_{1},x_{2},\ldots\}$ of elements of $S^{Y}$, a right-linear combination over $X$ is one of the form $(x_{1}\otimes\lambda_{1})\oplus(x_{2}\otimes\lambda_{2})\oplus\ldots$ for some $\lambda_{1},\lambda_{2},\ldots$ in $S$, such that only a finite number are different from $0_{S}$. The right-semimodule generated by $X$ is defined as all the elements of $S^{Y}$ that can be written as a right-linear combination over $X$. An element of the right-semimodule generated by $X$ will simply be said to be right-generated by $X$.
We will use the notation $[x]$ for elements of $S^{\Sigma^{*}}$; thus for a word $w\in\Sigma^{*}$ we write $[x]_{w}$ to denote the $w$-th component of $[x]$. The restriction of an element of $S^{\Sigma^{*}}$ to entries indexed by a subset $Z\subseteq\Sigma^{*}$ will be denoted by $[x]_{Z}$. Given an element $[x]\in S^{\Sigma^{*}}$, and a letter $a$ of $\Sigma$, we denote by $[ax]$ the element of $S^{\Sigma^{*}}$ defined by $[ax]_{w}=[x]_{wa}$ for all words $w$. If $x$ is a word and $f$ and $F$ are clear from context, we shall write $[x]$ to denote the element of $S^{\Sigma^{*}}$ indexed by $x$ in $F$. Note that if $x$ is a word, $f$ a function and $F$ its Hankel matrix, this definition is consistent: $[ax]_{u}=F_{u,ax}=f(uax)=F_{ua,x}=[x]_{ua}$. For a subset $X\subseteq S^{\Sigma^{*}}$ and $[x]$ in $X$, we will sometimes simplify notation by writing $x\in X$.
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) right-closed if for all $[x]$ in $X$ and $a\in\Sigma$, $[ax]$ is right-generated by $X$; (ii) right-closed on $Z$ if for all $[x]$ in $X$ and $a\in\Sigma$, $[ax]_{Z}$ is right-generated by the elements of $X$ restricted to $Z$; (iii) column-generating for $f$ if all the columns of the Hankel matrix of $f$ are right-generated by $X$. By an abuse of notation, we say that a set of words $Q$ is right-closed (resp. right-closed on $Z$, column-generating for $f$) if the set $X=\{[q]:q\in Q\}$ has this property.
Co-Hankel automata and co-hypothesis automata:
Notation A.1.
Given a function $f:\Sigma^{*}\rightarrow S$, a finite subset $T$ of $S^{\Sigma^{*}}$ and a subset $Q$ of $\Sigma^{*}$, we define $\Gamma_{Q,T}\subseteq S^{T}\times S^{T\times\Sigma\times T}$ to be the set of tuples $\lambda=(\lambda_{q}^{\varepsilon},\lambda_{q,a,p})_{q\in T,a\in\Sigma,p\in T}$ such that:
  • $[\varepsilon]_{Q}=\bigoplus_{q\in T}[q]_{Q}\otimes\lambda_{q}^{\varepsilon}$, and
  • for all $q\in T$ and all $a\in\Sigma$, $[aq]_{Q}=\bigoplus_{p\in T}[p]_{Q}\otimes\lambda_{q,a,p}$.
Note that $\Gamma_{Q,T}$ is non-empty if and only if $T$ is right-closed on $Q$ and $[\varepsilon]_{Q}$ is right-generated by $T$ restricted to $Q$. Note also that if $Q\subseteq Q’$ ’, then $\Gamma_{Q,T}\supseteq\Gamma_{Q^{\prime},T}$. If $Q=\Sigma^{*}$, we simply write $\Gamma_{T}$.
Definition A.2.
Given a finite subset $T$ of $S^{\Sigma^{*}}$ and a subset $Q$ of $\Sigma^{*}$, for each $\lambda\in\Gamma_{Q,T}$, we define a finite weighted automaton $\mathcal{H}^{co}_{Q,T,\lambda}$ with:
  • finite set of states $T$;
  • initial-state vector $([q]_{\varepsilon})_{q\in T}$;
  • final-state vector $(\lambda^{\varepsilon}_{q})_{q\in T}$; and
  • for all $q,p\in T$ and all $a\in\Sigma$, a transition from $p$ to $q$ labelled by $a$ with weight $\lambda_{q,a,p}$.
If $Q=\Sigma^{*}$, we say $\mathcal{H}^{co}_{Q,T,\lambda}$ is a co-Hankel automaton and write simply $\mathcal{H}^{co}_{T,\lambda}$ to reduce notation. If $T$ is a finite set of columns of the Hankel matrix and $Q$ is finite, we say that $\mathcal{H}^{co}_{Q,T,\lambda}$ is a co-hypothesis automaton, and view $T$ as a finite set of words by identifying with an appropriate index set.
Theorem A.3.
Let $f:\Sigma^{*}\rightarrow S$ be a function. There exists a finite set $T\subseteq S^{\Sigma^{*}}$ that is right-closed and right-generates $\langle\varepsilon\rangle$ if and only if $f$ is computed by a finite weighted automaton $\mathcal{A}$ over $S$. Specifically, if there exists a finite set $T\subseteq S^{\Sigma^{*}}$ that is right-closed and right-generates $\langle\varepsilon\rangle$, then $\mathcal{H}^{co}_{T,\lambda}$ computes $f$ for all $\lambda\in\Gamma_{T}\neq\emptyset$, whilst if $f$ is computed by a finite weighted automaton $\mathcal{A}$ over $S$, then the set of all $[q]$ where $q$ is a state of $\mathcal{A}$ and $[q]_{w}$ is the weight of $w$ to $q$ in $\mathcal{A}$ is right-closed and right-generates every column of the Hankel matrix of $f$.
(Weakly/strongly) co-guessable functions and chain conditions:
Definition A.4.
A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly co-guessable if there exist a finite set $T$ of columns of the Hankel matrix of $f$ such that $\Gamma_{T}$ is non-empty.
Definition A.5.
A function $f:\Sigma^{*}\rightarrow S$ is co-guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Gamma_{Q,T}=\Gamma_{T}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is co-guessable. Furthermore, we say that a co-guessable function $f$ is strongly co-guessable if for all finite sets $T\subseteq\Sigma^{*}$ with $\Gamma_{T}\neq\emptyset$ there exists a finite set $Q\subseteq\Sigma^{*}$ such that $(Q,T)$ witnesses that $f$ is co-guessable.
Definition A.6.
Given a function $f:\Sigma^{*}\to S$ and $F$ its Hankel matrix, $f$ satisfies the weak co-ascending chain condition if for all chains of right-semimodules $X_{0}\subseteq X_{1}\subseteq X_{2}\subseteq\ldots$ such that
  • each $X_{i}$ is generated by a finite set of columns of $F$;
  • each column of $F$ belongs to some $X_{i}$
there exists $n$ such that for all $m\geq n$, $X_{m}=X_{n}$.
Definition A.7.
Let $T$ be a finite subset of $\Sigma^{*}$. A total $T$-chain is an infinite descending chain of solution spaces $\Gamma_{Q_{0},T}\supsetneq\Gamma_{Q_{1},T}\supsetneq\Gamma_{Q_{2},T}\supsetneq\cdots$ with $Q_{i}\subseteq Q_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $Q_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is column-bound if there exists a finite set of words $T$ such that $\Gamma_{T}\neq\emptyset$ and there is no total $T-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly column-bound if for every finite set of words $T$ such that $\Gamma_{T}\neq\emptyset$ there is no total $T$-chain.
Appendix B Examples
Neither weakly guessable nor weakly co-guessable ($f_{1}$):
We will give examples of functions that are neither weakly guessable nor weakly co-guessable for $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, $\mathbb{R}_{\max}$, $\mathbb{N}$, $\mathbb{R}_{\geq 0}$ and $\mathcal{P}_{\rm fin}(\Sigma^{*})$. We give the proofs for weakly guessable. The proofs for weakly co-guessable are dual.
Let $S$ be one of $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, or $\mathbb{R}_{\max}$, and consider the function $f_{1}$ computed by the automaton depicted in Figure 1(b) and $F$ its Hankel matrix over $S$. Let $w=ua^{n}$ and $w^{\prime}=a^{n^{\prime}}v$ where $u,v\in\Sigma^{*}$ are such that $u$ does not end with an $a$ and $v$ does not start with an $a$. The corresponding entry of the Hankel matrix is given by $F_{w,w^{\prime}}=f(ww^{\prime})={\max}\{f_{1}(u),n+n^{\prime},f_{1}(v)\}$. We claim that $f_{1}$ is neither weakly guessable nor weakly co-guessable. Suppose for contradiction that $f_{1}$ is weakly guessable. By Theorem 3.5 there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the rows of the Hankel matrix indexed by $W$ are row-generating. Let us write $w_{i}=u_{i}a^{n_{i}}$ for $i=1,\ldots,k$, where $u_{i}$ does not end with an $a$ and $n_{i}\geq 0$, and choose a positive integer $N$ such that $N\geq{\max}_{i\in\{1,\ldots,k\}}\{f_{1}(u_{i}),n_{i}\}$. Consider the row of the Hankel matrix indexed by $w=a^{N+1}b$. Since the rows indexed by $W$ are assumed to be row-generating, we have: $\langle w\rangle=\bigoplus_{i=1}^{k}\lambda_{i}\otimes\langle w_{i}\rangle$ for some $\lambda_{i}\in S$. Now set $u=ba^{N}$ and $v=ba^{N+1}$. We have:
$f_{1}(wu)=F_{w,u}=\langle w\rangle_{u}$ $=$ ${\max}_{i}\{\lambda_{i}+\langle w_{i}\rangle_{u}\}$
$=$ ${\max}_{i}\{\lambda_{i}+F_{w_{i},u}\}$
$=$ ${\max}_{i}\{\lambda_{i}+\max(f_{1}(u_{i}),n_{i},N)\}$
$=$ ${\max}_{i}\{\lambda_{i}+N\}={\max}_{i}\{\lambda_{i}\}+N$
and, similarly, $f_{1}(wv)={\max}_{i}\{\lambda_{i}\}+N+1$. But this immediately gives a contradiction, since $f_{1}(wu)=f_{1}(a^{N+1}b^{2}a^{N})=N+1=f_{1}(a^{N+1}b^{2}a^{N+1})=f_{1}(wv)$.
Let $S$ be one of $\mathbb{N}$, $\mathbb{R}_{\geq 0}$ and consider the function $f^{\prime}_{1}$ computed by the automaton depicted in Figure 1(d) and $F$ its Hankel matrix. Suppose that $f^{\prime}_{1}$ is weakly guessable. Then there exist a finite set of words $Q=\{w_{1},w_{2},\ldots,w_{k}\}$ such that the rows indexed by $Q$ left-generate the rows of $F$. Without loss of generality, we may assume that $w_{1}=\varepsilon$ and that $w_{i}$ is a power of $a$ for $i=1,\ldots m\leq k$. Let $w=a^{n}$ for some positive integer $n$. Then by assumption, considering columns $\varepsilon$ and $a$ of $F$, there exist $\lambda_{1},\ldots,\lambda_{k}$ in $S$ such that:
$2^{n}-1=\sum_{i=1}^{k}\lambda_{i}f^{\prime}_{1}(w_{i})\mbox{ and }2^{n+1}-1=\sum_{i=1}^{k}\lambda_{i}f^{\prime}_{1}(w_{i}a),$
which gives $\sum_{i=1}^{k}\lambda_{i}f^{\prime}_{1}(w_{i}a)=1+\sum_{i=1}^{k}2\lambda_{i}f^{\prime}_{1}(w_{i}).$ Moreover, for each $w_{i}$ that is a power of $a$, we have $f^{\prime}_{1}(w_{i}a)=2f^{\prime}_{1}(w_{i})+1$, whilst if $w_{i}$ is not a power of $a$, we have $f^{\prime}_{1}(w_{i}a)=0=f^{\prime}(w_{i})$. Thus when restricting attention to columns $\{\varepsilon,a\}$ it suffices to consider only the $w_{i}$ with $i\leq m$, giving:
$\sum_{i=1}^{m}\lambda_{i}(2f^{\prime}_{1}(w_{i})+1)=1+\sum_{i=1}^{m}2\lambda_{i}f^{\prime}_{1}(w_{i})$
giving $\sum_{i=1}^{m}\lambda_{i}=1$.
Suppose now that $n$ is (strictly) larger than the lengths of all the $w_{i}$. We have proved that there are $\lambda_{i}\geq 0$ such that $2^{n}-1=\sum_{i=1}^{m}\lambda_{i}f^{\prime}_{1}(w_{i})$ and $\sum_{i=1}^{m}\lambda_{i}=1$. But $12^{n}-$ 1 is strictly greater than any of the $f^{\prime}_{1}(w_{i})$, so cannot be obtained as a convex combination of them.
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{1}$ computed by the automaton depicted in Figure 1(f) and $F$ its Hankel matrix. If a row $\langle w\rangle$ of $F$ is left-generated by a finite set of rows, this in particular means (by restricting attention to column $\varepsilon$) that $\{a^{|w|_{a}},b^{|w|_{b}}\}$ can be written as $\bigcup_{i=1}^{k}X_{i}\{a^{n_{i}},b^{m_{i}}\}$ for some finite sets $X_{i}$ and non-negative integers $n_{i},m_{i}$. This is only possibly if exactly one of the sets $X_{i}$ is non-empty, and for this value $i$ we have that $X_{i}=\{\varepsilon\}$ and $n_{i}=|w|_{a}$ and $m_{i}=|w|_{b}$. It follows from this that for all pairs of non-negative integers $(n,m)$ a row-generating set of words must contain at least one word with $|w|_{a}=n$ and $|w|_{b}=m$.
Weakly guessable but neither guessable nor weakly co-guessable ($f_{2}$).
Let $S=\mathbb{Z}_{\max}$ or $S=\mathbb{R}_{\max}$ and consider the function $f_{2}$ computed by the automaton depicted in Figure 1(a), but viewed as an automaton over $S$ and let $F$ denote its Hankel matrix.
First, $f_{2}$ is weakly guessable. Indeed, it is easy to see that one can transform the automaton given in the figure into an equivalent literal automaton, by pushing the weights from the first transitions to final weights.
Second, $f_{2}$ is not weakly co-guessable. For all words $w$, we have $f_{2}(aw)=|aw|_{a}$, and $f_{2}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{2}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating. Let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
Finally, $f_{2}$ is not guessable. Suppose for contradiction that $f_{2}$ is guessable. Then there exist finite sets $Q$, $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. Without loss of generality, we may assume that for a fixed constant $N$ (to be chosen below) we have $n>N$ where $n$ is the length of the longest word in $T$ (since for all $T^{\prime}\supseteq T$ we also have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}=\Lambda_{Q}\neq\emptyset$). Hence for all $\lambda$ in $\Lambda_{Q,T}$, we have that $\mathcal{H}_{Q,T,\lambda}$ computes $f_{2}$. If $Q$ only contains words starting with an $a$, then for all $q\in Q$, $\langle qa\rangle_{T}=\langle q\rangle+1$ and $\langle qb\rangle_{T}=\langle q\rangle$, and it is easy to see that a hypothesis automaton contructed from $Q$, $T$ and a $\lambda$ of this shape would not compute $f_{2}$. A symmetric argument holds if $Q$ only contains words starting with $b$. If $Q$ contains a word $q$ starting with $a$ and a word $q^{\prime}$ starting with $b$, then, we may write $\langle qa\rangle_{T}=\max(\langle q\rangle_{T}+1,\langle q^{\prime}\rangle_{T}-n)$, $\langle qb\rangle_{T}=\max(\langle q\rangle_{T},\langle q^{\prime}\rangle_{T}-n)$, $\langle q^{\prime}a\rangle_{T}=\max(\langle q\rangle_{T}-n,\langle q^{\prime}\rangle_{T})$ and $\langle q^{\prime}b\rangle_{T}=\max(\langle q\rangle_{T}-n,\langle q^{\prime}\rangle_{T}+1)$. It is then easy to see that there exists a hypothesis automaton constructed from sets $Q$, $T$ and a $\lambda$ using the values from the linear combinations above that would not compute $f_{2}$; it would for example misbehave on a word of the form $qa^{n}bb^{2n}$. Indeed, since $q$ starts with $a$ the target function computes $|q|_{a}+n$. If there is no run from any initial state to $q$ in the hypothesis automaton, then the automaton computes $\infty$- ∞ and we are done. Otherwise, there is a run of weight $x+2n$, where $x$ is the maximal weight of a run from an initial state to $q$. Taking $N=|q|_{a}-x$ we see that $x+2n>x+n+N=|q|_{a}+n=f_{2}(qa^{n}bb^{2n})$.
Weakly co-guessable but neither co-guessable nor weakly guessable ($\bar{f_{2}}$):
Since $\mathbb{Z}_{\max}$ and $\mathbb{R}_{\max}$ are commutative, and $f_{2}$ is weakly guessable but neither guessable nor weakly co-guessable, it follows that $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable).
Strongly guessable but not weakly co-guessable ($f_{3}$ and $f_{5}$):
We will give examples of functions that are strongly guessable but not weakly co-guessable for $\mathbb{N}_{\max}$, $\mathbb{N}$, $\mathcal{P}_{\rm fin}(\Sigma^{*})$ and $\mathbb{R}_{\geq 0}$. In Lemma 4.3, it is shown that all the functions from the first three semirings are strongly row-bound and hence, any weakly guessable function is strongly guessable. It is also the case for $\mathbb{R}_{\geq 0}$ by Lemma 4.4. To prove that the functions under consideration are strongly guessable, it is then enough to prove that they are weakly guessable, or equivalently that they are computed by a literal automaton.
For $S=\mathbb{N}_{\max}$, consider the function $f_{3}$ computed by the automaton depicted in Figure 1(a) and $F$ its Hankel matrix. It is easy to see that $f_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f_{3}(aw)=|aw|_{a}$, and $f_{3}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
For $S=\mathbb{N}$, consider the function $f^{\prime}_{3}$ computed by the automaton depicted in Figure 1(c) and $F$ its Hankel matrix. It is easy to see that $f^{\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime}_{3}(aw)=2^{|aw|_{a}}$, and $f^{\prime}_{3}(bw)=2^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=2^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=2$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{3}$ computed by the automaton depicted in Figure 1(e) and $F$ its Hankel matrix. It is easy to see that $f^{\prime\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime\prime}_{3}(aw)=a^{|aw|_{a}}$, and $f^{\prime\prime}_{3}(bw)=b^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=a^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=b^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=a^{N+1}$ and $F_{b,w}=b$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
Finally, we give an example of a function over $S=\mathbb{R}_{\geq 0}$ that is strongly guessable but not weakly co-guessable. Consider the automaton obtained from that in Figure 1(c) by replacing all weights labelled by $b$ by $1$. In a similar manner to the previous cases, it is easy to see that the function computed, which we denote by $f_{5}$, is computed by a literal automaton and hence is weakly guessable. For all words $w$, we have $f_{5}(aw)=2^{|aw|_{a}}$, and $f_{5}(bw)=1$. Taking $Q=T=\{\varepsilon,a,b\}$ it is straightforward to verify that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ (indeed, for each of $\langle\varepsilon\rangle,\langle qa\rangle,\langle qb\rangle$ where $q\in Q$ we find that there is exactly one way to write the given vector as a linear combination of the $\langle q^{\prime}\rangle$ with $q^{\prime}\in Q$). Hence $f_{5}$ is strongly guessable. Suppose then for contradiction that $f_{5}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\sum_{i}2^{|w_{i}|_{a}+1}$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
Strongly co-guessable but not weakly guessable ($\bar{f_{3}}$):
Since $\mathbb{N}_{\max}$ and $\mathbb{N}$ are commutative, and $f_{3}$ and $f^{\prime}_{3}$ are strongly guessable but not weakly co-guessable, then $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable (using once more the fact that a function over a commutative semiring is (weakly/strongly) guessable if and only if its mirror is (weakly/strongly) co-guessable).